# Risk-Neutral Pricing Model of Uniswap Liquidity Providing Position: A Stopping Time Approach

### v2

@@ Abstract
In this paper, we introduce a novel pricing model for Uniswap V3, built upon stochastic processes and the Martingale Stopping Theorem. This model innovatively frames the valuation of positions within Uniswap V3. We further conduct a numerical analysis and examine the sensitivities through Greek risk measures to elucidate the model’s implications. The results underscore the model’s significant academic contribution and its practical applicability for Uniswap liquidity providers (LPs), particularly in assessing risk exposure and guiding hedging strategies.
@@
Uniswap, pricing model, greeks, perpetual option, risk management, hedging strategies
@@ I Introduction
Uniswap[1], launched in November 2018, represents a decentralized, open-source protocol for cryptocurrency exchange. As the first automated market maker (AMM) to achieve widespread adoption, Uniswap enables users to conduct token swaps directly from their wallets via liquidity pools. In these pools, liquidity providers (LPs) earn fees by contributing their assets.
@@
The introduction of Uniswap V3 brought significant advancements, most notably the concept of concentrated liquidity. This feature allows LPs to allocate their capital to specific price ranges, enhancing capital efficiency and yielding higher returns while simultaneously mitigating price risk. Furthermore, Uniswap V3 incorporates multiple fee tiers, thereby providing LPs with flexibility in selecting their compensation levels in accordance with their risk tolerance.
With the ongoing expansion of decentralized finance (DeFi) adoption, the demands on LPs in Uniswap v3 have intensified. LPs are now required to meticulously balance a myriad of factors, including risk, returns, and volatility, necessitating a sophisticated understanding and strategic approach to liquidity provision.
II Related Work and Limitations
II-A Loss-Versus-Rebalancing (LVR)
LVR is a metric designed to quantify the impact of adverse selection on automated market maker (AMM) LPs[2]. LVR captures the costs incurred by LPs due to stale prices that better-informed arbitrageurs exploit. Unlike impermanent loss, which only considers the starting and ending points of a trade, LVR compares the performance of LPs to a rebalancing portfolio, highlighting additional costs associated with providing liquidity.
LVR incorporates volatility into its calculations, making it a useful tool for selecting the optimal liquidity pool. However, it remains primarily a risk assessment tool. It does not evaluate the returns for LPs when exiting positions at different prices, nor does it provide guidance on the optimal range for market-making activities.
II-B Guillaume Lambert
Guillaume Lambert focused on pricing within the Black-Scholes-Merton (BSM) framework in his work[3][4]. He viewed liquidity positions as perpetual options and attempted to replicate liquidity positions with covered calls. He also extends his model towards implied volatility and hedging. However, a limitation of Lambert’s model is the specification of an expiration time T, which inherently makes it an approximation rather than a precise solution. Additionally, Lambert explored the use of stochastic processes to calculate the probability of prices remaining within a position range, which has greatly inspired our research. We aim to build upon this foundation by starting from stochastic processes to initiate our study.
II-C Martingale and first hitting problem with two boundaries
Unlike traditional derivatives with specific expiration dates, Uniswap V3 contracts are perpetual, persisting until the price breaches predefined boundaries or the contract holder voluntarily ceases market-making. In scenarios where the price crosses these preset boundaries, this constitutes a classic First Passage Problem, for which Darling and Siegert provided a derivation of the stopping time distribution using the Laplace Transform in 1953 [5]. We have further discovered that the expected value of the payoff under the risk-neutral measure contains exclusively terms of the form $\tau^{n}\exp(-r\tau)$ which are associated with the stopping time $\tau$. Therefore, we employ the Laplace Transform and its implications to derive the present value when the holder does not elect to terminate prematurely. In situations where the contract holder actively ceases market-making, we draw inspiration from the pricing methodology for Perpetual American Puts described by Shreve (2004) [6] and use numerical optimization to determine the optimal stopping strategy.
III Derivation
III-A Summary of Model Derivation
@@
Uniswap V3 can be divided into two primary pricing components: the Liquidity Provider (LP) segment and the Rebate segment. The LP component is structured to accommodate either European or American option styles. Under the European construct, the Uniswap contract realizes its terminal value solely upon the price breaching the predetermined upper or lower thresholds. In contrast, the American framework allows the contract holder the flexibility to terminate the automated market-making process at any chosen point in time. The valuation of the Rebate segment hinges on the method selected by the contract holder for Rebate withdrawal. The upper limit scenario assumes the possibility of continuous Rebate extraction at the precise point of accrual. Meanwhile, the lower limit approach considers a singular, comprehensive withdrawal of all Rebates coinciding with the termination of the contract.
III-B Divide and Conquer
@@
III-B1 Liquidity Provider segment
The unit price of the LP segment within the Uniswap V3 Contract, constrained by $S_{H}>S_{0}>S_{L}$, can be expressed as:
@@ III-C European LP Pricing
III-D American LP Pricing
In Euro-like pricing, liquidity providers can only exit when the price reaches the boundary. We further propose an American-like pricing model that allows liquidity providers to exit at any time. As Uniswap V3 contracts are perpetual, the termination of the contract relies exclusively on price movements, independent of time $t$. Therefore, it can be assumed that there exist boundaries determined solely by price levels. This implies that exit strategies should be based only on price considerations, analogous to the derivation of the perpetual American put option in Shreve’s work[6]. So we have:
@@ III-E Rebate Pricing
$=C_{a}L_{q}V_{LP}(P_{t})(\cdot\mathbb{E}\left[\int_{0}^{\tau}e^{-rt}dt\right]$
$=C_{a}L_{q}V_{LP}(P_{t})\cdot\mathbb{E}\left[\frac{1-e^{-r\tau}}{r}\right]$
$=\frac{C_{a}L_{q}V_{LP}(P_{t})}{r}\cdot(1-\mathbb{E}\left[{e^{-r\tau}}\right])$
@@
$\mathbb{E}(V_{\text{fee}}(t))=\mathbb{E}(C_{a}L_{q}V_{LP}(P_{t})e^{-r\tau}\tau)=C_{a}L_{q}V_{LP}(P_{t})\mathbb{E}(e^{-r\tau}\tau)$
@@
Figure 2 provides a Monte Carlo simulation to model the price dynamics of an asset following a Geometric Brownian Motion (GBM). The initial asset price $S_{0}$ was set to 1, with volatility $\sigma=0.6$, drift $\mu=0$, risk-free rate $r=0.04$, and boundary levels $H=1.2$ (upper bound) and $L=0.8$ (lower bound). We ran the simulation 10,000 times, calculating the payoff at each simulation based on whether the price reached the upper or lower boundary, then discounting the payoff to obtain the expected value. The Monte Carlo estimates closely matched the analytical solution derived from our new pricing formula, confirming the accuracy of our approach.
The final distribution of the outcomes, represented by a histogram, exhibited a bimodal shape. This reflects the nature of the problem, where the asset price is either stopped at the upper bound $H$ or the lower bound $L$ due to the stopping conditions inherent in the Uniswap V3 liquidity model. The two peaks correspond to the outcomes where the price reaches these boundaries, capturing the distinctive characteristics of liquidity provision in a decentralized exchange framework. This bimodal distribution highlights the risk/reward trade-offs faced by liquidity providers in such a system, where extreme price movements lead to clear-cut outcomes based on boundary interactions.
III-F Formulation of Case-Specific Models
Based on the derivations from the foundational section, we have obtained the pricing formulas for the following four scenarios. For the pricing formula of the European LP option problem under a GBM process, assuming that transaction fees can only be collected upon exiting the position, the pricing formula is as follows:
@@ III-F Formulation of Case-Specific Models
$=LP_{H}e^{\mu^{\prime}b^{\prime}}\frac{\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$+LP_{L}e^{\mu^{\prime}a^{\prime}}\frac{\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$+C_{a}L_{q}V_{LP}(P_{t})$
$(\frac{(b-a)(e^{\mu^{\prime}a^{\prime}}\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}}))}{\sqrt{2r+\mu^{\prime 2}}\operatorname{th}((b-a)\sqrt{2r+\mu^{\prime 2}})\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$-\frac{e^{\mu^{\prime}a^{\prime}}b^{\prime}\operatorname{ch}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}a^{\prime}\operatorname{ch}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\sqrt{2r+\mu^{\prime 2}}\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})})$
@@
$V_{t}$ $=LP_{H}e^{\mu^{\prime}b^{\prime}}\frac{\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$+LP_{L}e^{\mu^{\prime}a^{\prime}}\frac{\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$+\frac{C_{a}L_{q}V_{LP}(P_{t})}{r}(1$
$-\frac{e^{\mu^{\prime}a^{\prime}}\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})})$
For the pricing formula of the American LP option problem, assuming that transaction fees can only be collected upon exiting the position, the pricing formula is as follows:
$V_{t}=LP_{L_{2}}e^{\mu^{\prime}d^{\prime}}\frac{\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$+LP_{L_{1}}e^{\mu^{\prime}c^{\prime}}\frac{\operatorname{sh}(bd^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}+C_{a}L_{q}V_{LP}(P_{t})($
$\frac{(d-c)(e^{\mu^{\prime}c^{\prime}}\operatorname{sh}(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}}))}{\sqrt{2r+\mu^{\prime 2}}\operatorname{th}((d-c)\sqrt{2r+\mu^{\prime 2}})\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$-\frac{e^{\mu^{\prime}c^{\prime}}d^{\prime}\operatorname{ch}(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}c^{\prime}\operatorname{ch}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\sqrt{2r+\mu^{\prime 2}}\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})})$
When we assume that transaction fees can be continuously collected during the lifetime of the position, the pricing formula for the American LP option problem becomes:
$V_{t}$ $=LP_{L_{2}}e^{\mu^{\prime}d^{\prime}}\frac{\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$+LP_{L_{1}}e^{\mu^{\prime}c^{\prime}}\frac{\operatorname{sh}(bd^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$+\frac{C_{a}L_{q}V_{LP}(P_{t})}{r}(1$
$-\frac{e^{\mu^{\prime}c^{\prime}}\operatorname{sh}(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})})$
IV Sensitivity Analysis

### v3

@@ Abstract
In this paper, we introduce a novel pricing model for Uniswap V3, built upon stochastic processes and the Martingale Stopping Theorem. This model innovatively frames the valuation of positions within Uniswap V3. We further conduct a numerical analysis and examine the sensitivities through Greek risk measures to elucidate the model’s implications. Additionally, we demonstrate the model’s practical application by construct the implied volatility of the model and compare it to other exist model. The results underscore the model’s significant academic contribution and its practical applicability for Uniswap liquidity providers (LPs), particularly in assessing risk exposure and guiding hedging strategies.
@@
Uniswap, pricing model, greeks, perpetual option, risk management, hedging strategies, implied volatility
@@ I Introduction
I-A Background and Motivation
Uniswap, launched in 2018, is one of the leading decentralized exchanges (DEXs) in the cryptocurrency market, enabling users to swap tokens directly from their wallets through liquidity pools, without relying on traditional intermediaries. As the first automated market maker (AMM)[1] to achieve widespread adoption, Uniswap has played a pivotal role in the growth of decentralized finance (DeFi).
@@ I-A Background and Motivation
The introduction of Uniswap V3 marked a significant advancement in decentralized exchange technology, most notably through the concept of concentrated liquidity. This feature allows LPs to allocate capital to specific price ranges, improving capital efficiency and enhancing potential returns while simultaneously reducing price risk.
With the ongoing growth of DeFi adoption, LPs now face a more complex landscape, balancing a range of factors such as risk, returns, and volatility. This requires a sophisticated understanding of liquidity provision strategies.
In this work, we treat Uniswap V3 as an exotic perpetual derivative and use the Black-Scholes framework, along with the Martingale Stopping Theorem, to derive a closed-form pricing formula. Additionally, we perform a Monte Carlo simulation and sensitivity analysis to evaluate the model’s performance. We also compare our model with other state-of-the-art models to assess its advantages and limitations.
Looking ahead, Uniswap V4 is expected to introduce even more advanced features that will further enhance the efficiency and flexibility of decentralized exchange mechanisms. Our model is designed to incorporate these new features of Uniswap V4, ensuring its adaptability to future developments.
I-B Related Work
Our work builds on and extends two broad strands of literature: (i) DeFi research on AMMs, and (ii) traditional financial engineering studies on market-making. Below, we discuss the most relevant works and highlight how our research contributes to these fields.
I-B1 Automated Market Makers
A significant advancement in AMM research is the concept of Loss-Versus-Rebalancing (LVR), introduced by [2] to quantify the adverse selection costs faced by LPs in AMM. LVR measures the performance gap between an LP’s portfolio and a dynamically rebalanced portfolio, capturing the losses incurred due to stale prices exploited by arbitrageurs.
Another important contribution is the work of [3], who modeled Uniswap V3 liquidity positions as perpetual options within the Black-Scholes-Merton (BSM) framework. Lambert’s work treats LP positions as covered calls and explores implied volatility and hedging strategies.
Recent studies have further advanced the understanding of AMMs. For example, [4] analyze the tradeoff between risks and returns for LPs in Uniswap V3, while [5] study the predictable losses of LPs in a continuous-time. [6] explore the hedging of impermanent losses. Closest to our work are the models in [7] and [8], which focus on fee revenue and use approximation techniques to obtain dynamic strategies.
I-B2 Market-Making in Traditional Finance
Our work is also related to the market-making literature in traditional finance. Early works include [9] and [10], which model market-making as an inventory management problem and derive optimal bid-ask spread strategies. These foundational studies have been extended in many directions; see [11] and [12] for recent developments
Another particularly relevant area is the study of perpetual derivatives, which share similarities with the perpetual nature of Uniswap V3 liquidity positions. [13] provides a comprehensive treatment of perpetual put options using stochastic processes, deriving the distribution of stopping times and optimal stopping strategies.
I-C Contributions
Our work is structured into four key steps, each addressing a critical aspect of pricing Uniswap V3 liquidity positions and extending the model’s applicability:
Step 1: Perpetual Liquidity Position Modeling
We model Uniswap V3 liquidity positions as perpetual derivatives, using the Martingale Stopping Theorem and Laplace Transform to derive closed-form pricing formulas. This approach captures the perpetual nature of liquidity positions and eliminates the need for fixed expiration times.
Step 2: Monte Carlo Validation
We validate our closed-form formulas through Monte Carlo simulations. The results align closely with the analytical solutions, confirming the accuracy and robustness of our model.
Step 3: Greeks and Implied Volatility
We conduct a comprehensive analysis of Greek risk measures (Delta, Gamma, Vega) and derive implied volatility from our model. A key comparison with the LVR framework highlights the advantages of our approach in capturing implied volatility.
Step 4: Generalization to Other AMMs and Uniswap V4
We discuss the potential extension of our model to other AMMs and its adaptation to the dynamic fee structures expected in Uniswap V4. While our current focus is on Uniswap V3, we outline possible modifications to the framework to accommodate future developments in AMM design.
II Derivation
II-A Summary of Model Derivation
@@
Uniswap V3 can be divided into two primary pricing components: the LPs segment and the Rebate segment. The LPs component is structured to accommodate either European or American option styles. Under the European construct, the Uniswap contract realizes its terminal value solely upon the price breaching the predetermined upper or lower thresholds. In contrast, the American framework allows the contract holder the flexibility to terminate the automated market-making process at any chosen point in time. The valuation of the Rebate segment hinges on the method selected by the contract holder for Rebate withdrawal. The upper limit scenario assumes the possibility of continuous Rebate extraction at the precise point of accrual. Meanwhile, the lower limit approach considers a singular, comprehensive withdrawal of all Rebates coinciding with the termination of the contract.
II-B Divide and Conquer
@@
II-B1 Liquidity Provider segment
The unit price of the LPs segment within the Uniswap V3 Contract, constrained by $S_{H}>S_{0}>S_{L}$, can be expressed as:
@@ II-C European LP Pricing
II-D American LP Pricing
In Euro-like pricing, liquidity providers can only exit when the price reaches the boundary. We further propose an American-like pricing model that allows liquidity providers to exit at any time. As Uniswap V3 contracts are perpetual, the termination of the contract relies exclusively on price movements, independent of time $t$. Therefore, it can be assumed that there exist boundaries determined solely by price levels. This implies that exit strategies should be based only on price considerations, analogous to the derivation of the perpetual American put option in [13]’s work. So we have:
@@ II-E Rebate Pricing
[…]
$=C_{a}L_{q}\cdot\mathbb{E}\left[\frac{1-e^{-r\tau}}{r}\right]$
$=\frac{C_{a}L_{q}}{r}\cdot(1-\mathbb{E}\left[{e^{-r\tau}}\right])$
@@
$\mathbb{E}(V_{\text{fee}}(t))=\mathbb{E}(C_{a}L_{q}e^{-r\tau}\tau)=C_{a}L_{q}\mathbb{E}(e^{-r\tau}\tau)$
@@
Figure 2 shows a Monte Carlo simulation of asset price dynamics under GBM, with $S_{0}=1,\sigma=0.6,\mu=0,r=0.04,H=1.2,L=0.8$.Running 10,000 simulations, we calculated payoffs based on boundary hits, discounted them, and matched the results to our analytical solution, confirming accuracy.
The outcome histogram exhibits a bimodal shape, reflecting price stops at $H$ or $L$ due to Uniswap V3’s stopping conditions. The peaks correspond to boundary interactions, capturing the risk/reward trade-offs for liquidity providers in decentralized exchanges.
II-F Formulation of Case-Specific Models
Based on the derivations from the foundational section, we have obtained the pricing formulas for the following four scenarios. For the pricing formula of the European LPs option problem under a GBM process, assuming that transaction fees can only be collected upon exiting the position, the pricing formula is as follows:
@@ II-F Formulation of Case-Specific Models
$=LP_{H}e^{\mu^{\prime}b^{\prime}}\frac{sh(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$+LP_{L}e^{\mu^{\prime}a^{\prime}}\frac{sh(b^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$+C_{a}L_{q}*$
$(\frac{(b-a)(e^{\mu^{\prime}a^{\prime}}sh(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}sh(a^{\prime}\sqrt{2r+\mu^{\prime 2}}))}{\sqrt{2r+\mu^{\prime 2}}th((b-a)\sqrt{2r+\mu^{\prime 2}})sh((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$-\frac{e^{\mu^{\prime}a^{\prime}}b^{\prime}ch(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}a^{\prime}ch(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\sqrt{2r+\mu^{\prime 2}}sh((b-a)\sqrt{2r+\mu^{\prime 2}})})$
@@
$V_{t}$ $=LP_{H}e^{\mu^{\prime}b^{\prime}}\frac{sh(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$+LP_{L}e^{\mu^{\prime}a^{\prime}}\frac{sh(b^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$+\frac{C_{a}L_{q}}{r}(1$
$-\frac{e^{\mu^{\prime}a^{\prime}}sh(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}sh(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((b-a)\sqrt{2r+\mu^{\prime 2}})})$
For the pricing formula of the American LPs option problem, assuming that transaction fees can only be collected upon exiting the position, the pricing formula is as follows:
$V_{t}=LP_{L_{2}}e^{\mu^{\prime}d^{\prime}}\frac{sh(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$+LP_{L_{1}}e^{\mu^{\prime}c^{\prime}}\frac{sh(bd^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$+C_{a}L_{q}(\frac{(d-c)(e^{\mu^{\prime}c^{\prime}}sh(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}sh(c^{\prime}\sqrt{2r+\mu^{\prime 2}}))}{\sqrt{2r+\mu^{\prime 2}}th((d-c)\sqrt{2r+\mu^{\prime 2}})sh((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$-\frac{e^{\mu^{\prime}c^{\prime}}d^{\prime}ch(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}c^{\prime}ch(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\sqrt{2r+\mu^{\prime 2}}sh((d-c)\sqrt{2r+\mu^{\prime 2}})})$
When we assume that transaction fees can be continuously collected during the lifetime of the position, the pricing formula for the American LPs option problem becomes:
$V_{t}$ $=LP_{L_{2}}e^{\mu^{\prime}d^{\prime}}\frac{sh(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$+LP_{L_{1}}e^{\mu^{\prime}c^{\prime}}\frac{sh(bd^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$+\frac{C_{a}L_{q}}{r}(1$
$-\frac{e^{\mu^{\prime}c^{\prime}}sh(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}sh(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((d-c)\sqrt{2r+\mu^{\prime 2}})})$
III Sensitivity Analysis